0850: Day starts: Drank 1L of water to hydrate body and counteract dehydration caused by long term exposure to dry, cool air conditioned laboratory.
0855: System checks. Inspected the ion gauge pressures, the status of all pumps and LN2 level. Inserted MIG into beam path in preparation for BEP flux measurements.
0856: Outgas Ga. Executed a program in the Ga cell’s PID controller than ramps up to 975°C in 30 minutes, holds for 30 minutes and then ramps down to 950°C and holds.
0900: Sample load/unload. The two samples in the FEL (grown yesterday) were removed and placed in labelled boxes. Two new Si(111) substrates were mounted and the FEL pumped down manually and the degas cycle started . The intention to implement “automatic FEL pump down and degas” was reaffirmed whilst shaking a fist at the heavens. Samples were passed to a colleague for SEM analysis.
0920: Update log. Sample locations were updated on the “locations” tab of the log spread sheet to reflect new sample locations and status.
0930: As flux check. As shutter and valve were opened. Fluxes were taken at several valve positions consistent with previous days. Fluxes were recorded in the “flux check record” tab of the log spread sheet.
0945: As flux check completed. Gathered flux values were compared with legacy data for continuity checks. As shutter was closed and background As4 was pumped from the system. Vacuum recovery progressed prior to Ga flux check.
0950: Ga flux check. Ga shutter was opened for 60 seconds and the flux was recorded, averaged over 10 seconds. The shutter was then closed, the background flux was sampled, subtracted and the “flux check record” tab of the log spread sheet was updated. Current flux value was compared to legacy data for continuity.
1000: Sample 1: Start. Transferred sample from outgas stage to growth chamber and manually began recipe. Sample growth sheet was created.
1020: Confer with colleague. Discussed recent results with a colleague and outline strategy for samples to grow today.
1110: Platen cleaning P50 dry abrasive sheet. Cleaned worked ok but glue was water soluble and grit came away from cloth. Will need to order wet and dry P60 ASAP.
1145: Sample 2 Outgas. Transferred sample2 from FEL to outgas stage and executed outgas program in PID controller.
1200: Sample 1: Complete. Transferred sample from growth chamber to FEL. Transferred sample 2 from outgas stage to growth chamber. Transferred new sample from FEL to outgas stage and executed outgas program in PID controller.
1230: Sample 2: Start. Manually began sample 2 recipe. Initial step is oxide conditioning, which takes 45 minutes including the temperature ramp from 100°C to 700°C. Sample growth sheet was created.
1240: Lunch and consumption of another litre of water.
1530 Sample 2: End. After cooling, sample 2 was transferred from the growth chamber to the FEL.
1600 Sample load/unload. Both sample 1 and sample 2 were removed and placed in labelled boxes. Two new Si(111) substrates were mounted and the FEL pumped down manually and the degas cycle started . The intention to implement “automatic FEL pump down and degas” was reaffirmed whilst shaking a fist at the heavens. Samples were passed to a colleague for SEM analysis.
1615: Paperwork. Growth sheets were double checked. Run log was updated. All files were backed up to online server. Paper lab book was updated.
1700: Day ends. Drink another litre of water. Go home.
Total 8 hours. Growth: 2 samples. Samples/hour: 0.25.